Question 3: Suppose the inverse demand function for good X is given as P=100-2Q. Find

the price elasticity of demand when Q=40 units. What can you say about the type of the good

looking at the price elasticity of demand that you calculate in the first part?


1
Expert's answer
2022-01-31T10:12:29-0500

"P=100-2Q\\
Solve for Q\\
Q=\frac{100-P}{2}\\
Q= 50-\frac{P}{2}\\
Q=40units\\
40=50-\frac{P}{2}\\
P=20\\
E_p=\frac{dQ}{dP}×\frac{P}{Q}\\
Where \frac{dQ}{dP}= slope=50\\
E_p=50×\frac{20}{40}=25\\"

The demand is elastic; i.e the quantity changes faster than price. 1% increase in price= 25% increase in the quantities demanded.